OpenStudy (anonymous):

What is the simplified form of x+5/x^2-25 * x+5/x^2+10x+25?

OpenStudy (anonymous):

(x+5)/(x^2-25) = (x+5)/((x+5)(x-5)) = 1/(x-5) So (x+5)/(x^2-25) = 1/(x-5) ------------------------------------------------------- (x-5)/(x^2-10x+25) = (x-5)/((x-5)(x-5)) = 1/(x-5) So (x-5)/(x^2-10x+25) = 1/(x-5) so the original would turn to (1/(x-5))*(1/(x-5)) which then multiplies out to 1/((x-5)(x-5)) 1/(x^2-10x+25)

OpenStudy (anonymous):

get it?